Inlet ‘1’ and ‘2’ together can fill a tank in 18 hours. Inlet ‘2’ can fill the tank on its own in 24 hours. In how many hours can inlet ‘1’ alone fill the tank?
Answer by nerdybill(7384)

Inlet ‘1’ and ‘2’ together can fill a tank in 18 hours. Inlet ‘2’ can fill the tank on its own in 24 hours. In how many hours can inlet ‘1’ alone fill the tank?
.
Let x = time (hours) for inlet 1 alone
then
18(1/x + 1/24) = 1
multiplying both sides by 24x:
18(24 + x) = 24x
432 + 18x = 24x
432 = 6x
72 hours = x